# Urban Region Pre-training and Prompting: A Graph-based Approach

### v1

@@ 4.3. Multi-view Representation Learning
We utilize a triplet network (Hoffer and Ailon, 2015) to capture region similarity based on spatial proximity. For a given subgraph $g_{r}$ of region $r$, which serves as the anchor sample, we construct its positive and negative samples based on spatial proximity. If region $r^{\prime}$ is adjacent to $r$ (i.e.,  there is a “NearBy” type of edge connecting the corresponding nodes), we take the subgraph of $r^{\prime}$ as the positive sample, denoted by $g_{r}^{+}$. Otherwise, we take the subgraph as the negative sample, denoted by $g_{r}^{-}$. We input $g_{r}$, $g_{r}^{+}$, and $g_{r}^{-}$ to the graph encoder $\textsf{gEnc}(\cdot)$, obtaining the vector representations $\vec{h}_{r}$, $\vec{h}_{r}^{+}$, and $\vec{h}_{r}^{-}\in\mathbb{R}^{d}$, respectively. Then we use triplet loss to minimize the distance between adjacent regions and maximize it between distant regions, with the loss function calculated:
@@
where $\delta$ represents the margin parameter of triplet loss.
@@
Urban imagery encompasses various types of image information that describe the visual characteristics of a city, such as satellite images, street view images, and more. In this paper, we use satellite images as the source of urban imagery. For a given region $r$, its imagery can be represented as the set $\{a_{1},a_{2},\cdots,a_{m}\}$. We employ a pre-trained visual model to extract visual features of the region $\vec{h}_{r}^{img}$ and fuse these features with the region’s representation with the contrastive loss $\mathcal{L}_{img}$ (see Appendix A.2 for more details).
4.3.3. Human mobility.
Human mobility describes the dynamics of urban systems and is typically represented by trajectory data. For a given region $r$, the outflow feature $\mathbf{s}_{r}$ and inflow feature $\mathbf{t}_{r}$ are defined by calculating the number of outgoing and incoming trips within each time period. This results in the outflow and inflow feature sets $\{\mathbf{s}_{1},\mathbf{s}_{2},\cdots,\mathbf{s}_{N}\}$ and $\{\mathbf{t}_{1},\mathbf{t}_{2},\cdots,\mathbf{t}_{N}\}$, respectively, where $\mathbf{s}_{i},\mathbf{t}_{i}\in\mathbb{R}^{l}$, and $l$ is the number of intervals (e.g., 24). We then encode the inflow and outflow features using an encoder (such as a Multi-Layer Perceptron) to obtain $d$-dimensional embeddings:
$\vec{h}^{\text{src}}_{r}=\textsf{Encoder}(\mathbf{s}_{r})\quad,\quad\vec{h}^{\text{dst}}_{r}=\textsf{Encoder}(\mathbf{t}_{r}).$
Here we adopt the reconstruction loss $\mathcal{L}_{flow}$ to optimize the two embeddings (see Appendix A.3 for more details).
4.3.4. Region representation learning.
@@
Here, $\vec{h}^{k}_{r}$ ($k=1,2,3,4$) represents one of the four embeddings, and $\textsf{Decoder}^{k}$ is the corresponding decoder (such as a Multi-Layer Perceptron).
The final loss function can be formulated as follows:
@@
5. Prompt Task-specific Knowledge
After obtaining a general region representation, a critical issue lies in how to integrate task-specific knowledge into pre-trained representations. To achieve this, we propose two graph-based prompt tuning methods: a manually-designed prompt method based on explicit domain knowledge and a task-learnable prompt learning method to extract implicit task-specific knowledge.
@@ 5.1. Manually-designed Prompt
In the manually-designed prompt, we adjust the subgraph structure, i.e., , defining the graph pattern $\mathcal{P}(P_{V},P_{E},\theta)$ in Section 4.2, to emphasize specific relationships and entities relevant to the target task, as shown in Figure 5.
We further enhance this process by manually adjusting the proportion of different types of edges and entities in the subgraph, thereby influencing the resultant embeddings. To achieve this, we first define a set of task-specific entity weights $\mathbf{W}_{task}$ that reflect the importance of various types of entities for the downstream task. These weights are designed based on domain knowledge and the specific requirements of the task. Formally, we define $\mathbf{W}_{task}$ as:
$\mathbf{W}_{task}=\{w_{\tau_{i}}\mid\tau_{i}\in P_{V}\}.$
where $P_{V}$ is the set of all possible node types in the region-induced subgraph, and $w_{\tau_{i}}\in[0,1]$ represents the importance of entity type $\tau_{i}$. The weights $w_{\tau_{i}}$ are manually defined based on domain knowledge and are used as the deletion proportion for each entity type. We define a function $\textsf{Adjust}(\cdot)$ to delete nodes based on $\mathbf{W}_{task}$$\textsf{Adjust}(g_{r},\mathbf{W}_{task})\rightarrow g^{\prime}_{r}$ where $g_{r}=(V_{r},E_{r})$ is the original subgraph. The function Adjust obtains the adjusted subgraph $g^{\prime}_{r}=(V^{\prime}_{r},E^{\prime}_{r})$ as follows:
@@
The adjusted subgraph $g^{\prime}_{r}=(V^{\prime}_{r},E^{\prime}_{r})$ is then input to the pre-trained GURP to generate the task-specific embeddings.
In summary, the manually-designed prompt is the graph pattern $\mathcal{P}$ with the set of task-specific entity weights $\mathbf{W}_{task}$, which guide the $\textsf{Adjust}(\cdot)$ function in modifying the subgraph $g_{r}$ to $g^{\prime}_{r}$. By incorporating these manually-designed prompts, we can tailor the urban region representations to better capture task-relevant entities and relationships. This approach also offers the advantage of adapting to downstream tasks without the need for additional training. It simply requires adjusting the input subgraph based on the manually-defined prompts.
Figure 5. Manually-designed Prompt.
@@ 5.2. Task-learnable Prompt
Apart from leveraging explicitly designed task-specific knowledge prompts to enrich region representations, in the presence of sufficient task data, we can also learn prompts based on the task data and pre-trained representations to induce implicit knowledge within the task data, thereby enhancing region representations.
Considering that we describe the attributes of an urban region through subgraph induced with region nodes from $G$, it is intuitive to similarly describe the task adaptation to different region attributes with graphs as well. Therefore, we propose a prompt learning module to use prompt graphs to portray the adaptation of downstream task relative to the original region embedding, as shown in Figure 6.
@@
We introduce the graph kernel method (Vishwanathan et al., 2010; Nikolentzos and Vazirgiannis, 2020; Feng et al., 2022) to assess the similarity of between the prompt graphs and the region subgraphs due to its excellent performance and computational efficiency, then enhancing the pre-trained region embedding. Let $\mathcal{G}$ denote the prompt graph set consisting of $m$ trainable (hidden) prompt graphs, where each prompt graph $G^{p}\in\mathcal{G}$ is defined as $G^{p}=(V^{p},E^{p},\mathbf{X}^{p})$. Here, $\mathbf{X}^{p}\in\mathbb{R}^{|V^{p}|\times d}$ represents the node attributes, with $d$ being the dimension of these attributes. For each region subgraph $g_{r}$, we utilize an attribute graph kernel function $K_{p}(\cdot,\cdot)$ to measure the similarity between $g_{r}$ and each prompt graph $G^{p}_{1},G^{p}_{2},\ldots,G^{p}_{m}$. Specifically, we employ the $P$-step random walk kernel (more details in Appendix A.4) to compute the kernel value between the region subgraph $g_{r}$ and each prompt graph $G^{p}_{i}$. These values are concatenated to form an $m$-dimensional prompt vector $\vec{h}_{p}\in\mathbb{R}^{m}$ for each region. We concatenate the prompt vector $\vec{h}_{p}\in\mathbb{R}^{m}$ before the pre-trained region representation $\vec{h}_{r}$ to derive region embedding $\vec{h}^{\prime}_{r}=\vec{h}_{p}||\vec{h}_{r}$. The new embeddings are then fed into a linear layer $\textsf{Linear}(\cdot)$. We update the prompt graphs and linear layer parameters, by optimizing on the downstream labeled task, to obtain the final prompt representation with the simple MSE loss function:
$\mathcal{L}_{task}=\frac{1}{N}\sum_{i=1}^{N}(y^{pred}_{i}-y^{label}_{i})^{2}.$
@@ 6. Experiments
We conduct extensive experiments to investigate the following research questions (RQs):
  • RQ1: Can our proposed pre-train and prompt framework (GURPP) outperform previous approaches?
  • RQ2: Does GURP have the ability to capture and apply heterogeneous and generic patterns?
  • RQ3: How does the manually-designed prompt work?
  • RQ4: How does the task-learnable prompt perform and how does it look like?
  • RQ5: How does each component contribute to GURPP?
@@ 6.1. Experimental Setup
We collect the datasets of region division, POI, urban imagery, taxi trips, road network, check-in, crime and crash for two representative cities in the United States: New York City (NYC) and Chicago (CHI). The detailed statistics and sources of the dataset is available in Appendix A.5.
@@
We compare our model with the following representative baseline models: knowledge graph embedding methods and its variants (TransR (Lin et al., 2015), TransR-N, TransR-G), graph embedding methods (node2vec (Grover and Leskovec, 2016), GAE (Kipf and Welling, 2016)), and urban region embedding methods (MVURE (Zhang et al., 2021), MGFN (Wu et al., 2022), HREP (Zhou et al., 2023a), HRE, ReCP (Li et al., 2024)). These baselines cover the range from traditional graph embedding to complex models that take into account urban-specific attributes, providing us with a comprehensive performance comparison benchmark. Specific details and descriptions can be found in the Appendix A.6.
@@
To evaluate the prediction performance, we use three widely recognized evaluation metrics: the coefficient of determination ($R^{2}$), root mean squared error (RMSE), and mean absolute error (MAE). Better performance is indicated by a higher $R^{2}$ and lower RMSE and MAE values. Our model includes three version: 1) GURP represents the pre-train module, 2) $\text{GURPP}_{M}$ is GURP with manually-designed prompt, and 3) $\text{GURPP}_{T}$ is GURP with task-learnable prompt. For detailed implementation, please refer to Appendix A.7 and our provided code repository.
Table 1. Main performance comparison of different models on downstream tasks across NYC and CHI datasets. The best results for each metric are highlighted in bold, the second-best are underlined, and the third-best are marked with an asterisk. The last and the third-to-last rows show the percentage improvement of the GURPP and GURP models over the third-best model.
Model New York City (NYC) Chicago (CHI)
Crime Prediction Check-in Prediction Crime Prediction Crash Prediction
MAE RSME $R^{2}$ MAE RSME $R^{2}$ MAE RSME $R^{2}$ MAE RSME $R^{2}$
TransR-N 111.319 148.863 -0.014 534.699 805.282 0.031 209.072 307.454 -0.037 106.914 176.860 0.001
TransR-G 109.221 146.905 0.013 530.427 802.426 0.038 208.404 306.155 -0.028 104.680 174.419 0.028
node2vec 138.458 175.797 -0.414 573.118 837.441 -0.048 246.066 349.965 -0.343 121.900 174.907 0.023
GAE 96.846 126.364 0.270 349.942 542.920 0.560 200.606* 300.237 0.011 100.536 173.323 0.040
MVURE 72.471* 99.844 0.544 310.973 498.076 0.629 206.080 289.859* 0.079* 96.953 148.865* 0.292*
MGFN 73.502 97.440* 0.566* 300.150 456.626* 0.688* 214.021 310.832 -0.059 108.3740 179.854 -0.033
HRE 76.926 102.07 0.523 329.547 513.801 0.606 206.612 298.389 0.024 99.952 158.625 0.196
HREP 74.854 100.035 0.542 320.335 505.516 0.618 206.591 298.238 0.025 99.766 158.658 0.196
ReCP 73.853 100.558 0.537 292.871* 524.890 0.588 207.485 298.965 0.020 93.298* 157.515 0.207
GURP 72.094 93.216 0.603 271.689 412.918 0.745 161.266 225.244 0.444 80.362 108.781 0.622
Enhance 0.52% 4.33% 6.54% 7.23% 9.57% 8.28% 19.61% 22.29% 462.03% 13.87% 26.93% 113.01%
GURPP 71.481 91.612 0.616 261.450 388.225 0.775 146.385 203.050 0.548 65.297 92.114 0.729
Enhance 1.36% 5.98% 8.83% 10.73% 14.98% 12.65% 27.03% 29.95% 593.67% 30.01% 38.12% 149.66%
6.2. RQ1 & RQ2: Overall Performance
@@
We evaluate the overall performance of baselines and our model on three downstream tasks. The results of the experiments are shown in Table 1, where the GURPP is all implemented by $\text{GURPP}_{T}$. From the results, we find the following:
(1) GURPP demonstrates superior performance to baselines in both pre-training and prompting phases. In the pre-training phase, the performance of GURP improves by 6.54%, 8.28%, 462.03%, and 113.01% over the best baseline of $R^{2}$ on two urban datasets. Meanwhile, on all downstream tasks, the average performance on MAE and RMSE metrics improves by 17.28% and 22.26%, respectively. Subsequently, the performance of GURPP is further improved by an average of 9.50%, 9.28%, and 11.703% of MAE, RMSE and $R^{2}$, respectively, when we introduce the task-learnable prompt. The results validate the significant advantages of our proposed pre-train and prompt framework.
@@
(2) To validate the ability of our pre-training method to capture generic urban knowledge, we compare the performance of state-of-the-art urban region embedding methods on datasets from two distinct cities, NYC and CHI. The results show that while these baselines perform well on the NYC dataset, their performance drops significantly on the CHI dataset. For instance, the performance on CHI decreases by an average of 71.8% in $R^{2}$ compared to NYC. In contrast, our GURPP model demonstrates remarkable consistency, with only an 8.49% decrease in $R^{2}$. This may be due to the fact that the baselines tend to be effective only at extracting information specific to certain data types, but they do not consider spatial structure and functional layouts among urban entities. For example, existing baselines rely on the flow view to extract information. However, the content of flow view in the CHI dataset is sparse, leading to a significant degradation in the performance of downstream task. While GURPP shows excellent performance on both urban datasets, which demonstrates the robustness and universality of GURPP across diverse urban environments.
@@
(3) To further validate the effectiveness of our subgraph-centric and multi-view fusion module, we conduct experiments comparing the performance of models with and without subgraphs and multi-view fusion. Specifically, we compare the TransR-G model (adopts region-induced subgraph and regards graph-level embedding as the region embedding) with the TransR-N model (directly uses region nodes embedding as region embedding). The results indicate that TransR-G outperforms TransR-N by an average of 1.27% in MAE, underscoring the effectiveness of the subgraph-centered approach. This improvement is attributed to the subgraph-centric method’s ability to better preserve the structural and contextual information of regions, providing more accurate and meaningful urban region embeddings. The average 32.55% improvement in MAE for GURP over TransR-G is mainly due to the multi-view fusion strategy, which improves the model’s comprehensive understanding of urban features.
6.3. RQ3: Manually-designed Prompt Study
In the real world, the incidence of crime is often closely linked to specific POI categories, roads and junctions, for example, ATMs, bars, parks, etc. are more likely to be the location of a crime. The density of roads and the complexity of junctions are also important factors affecting crime rates. As for the check-in behavior, the brand of the POI is given more attention, and certain users may be more inclined to check in at locations of well-known brands such as Starbucks, McDonald’s, or other specific brands. Therefore, we designed two patterns as shown in Figure 7 to validate the performance of manually-designed prompt.
The results are shown in Table 2, from which it can be found that crime prediction does pay more attention to factors such as POI categories, roads and junctions, which are closely related to criminal behavior. Meanwhile, removing POI brand information shows a significant decrease in the performance of check-in prediction. These results also demonstrate that the prediction performance of the model can be improved by manually-designed prompt.
Figure 7. The structure of the manually-designed prompt based on Pattern 1 (P1) and Pattern 2 (P2). P1 ignores POI brand information. P2 randomly samples POI, road, and junction nodes around the region with a weight of 0.9.
Table 2. Performance of $\text{GURPP}_{M}$.

### v4

@@ 3.3. Multi-view Self-supervised Learning
We utilize a triplet network (Hoffer and Ailon, 2015) to capture region similarity based on spatial proximity. For a given subgraph $g_{r}$ of region $r$, which serves as the anchor sample, we construct its positive and negative samples based on spatial proximity. If region $r^{\prime}$ is adjacent to $r$ (i.e.,  there is a “NearBy” type of edge connecting the corresponding nodes), we take the subgraph of $r^{\prime}$ as the positive sample, denoted by $g_{r}^{+}$. Otherwise, we take the subgraph as the negative sample, denoted by $g_{r}^{-}$. We input $g_{r}$, $g_{r}^{+}$, and $g_{r}^{-}$ to the graph encoder $\textsf{gEnc}(\cdot)$, obtaining the vector representations $\vec{h}_{r}$, $\vec{h}_{r}^{+}$, and $\vec{h}_{r}^{-}\in\mathbb{R}^{d}$, respectively. Then we use triplet loss to minimize the distance between adjacent regions and maximize it between distant regions, where $\delta$ represents the margin parameter of triplet loss:
@@
Urban imagery encompasses various types of image information that describe the visual characteristics of a city, such as satellite images, street view images, and more. In this paper, we use satellite images as the source of urban imagery. For a given region $r$, its imagery can be represented as the set $\{a_{1},a_{2},\cdots,a_{m}\}$. We employ a pre-trained visual model to extract visual features of the region $\vec{h}_{r}^{img}$ and fuse these features with the region’s representation with the contrastive loss $\mathcal{L}_{img}$ (see Appendix A.2 for details).
3.3.3. Mobility pattern.
Mobility pattern describes the dynamics of urban systems and is typically represented by trajectory data. For a given region $r$, the outflow feature $\mathbf{s}_{r}$ and inflow feature $\mathbf{t}_{r}$ are defined by calculating the number of outgoing and incoming trips within each time period. This results in the outflow and inflow feature sets $\{\mathbf{s}_{1},\mathbf{s}_{2},\cdots,\mathbf{s}_{N}\}$ and $\{\mathbf{t}_{1},\mathbf{t}_{2},\cdots,\mathbf{t}_{N}\}$, respectively, where $\mathbf{s}_{i},\mathbf{t}_{i}\in\mathbb{R}^{l}$, and $l$ is the number of intervals (e.g., 24). We encode the inflow and outflow features with an encoder (such as a Multi-Layer Perceptron) to obtain $d$-dimensional embeddings: $\vec{h}^{\text{src}}_{r}=\textsf{Encoder}(\mathbf{s}_{r})$ and $\vec{h}^{\text{dst}}_{r}=\textsf{Encoder}(\mathbf{t}_{r}).$ Here we adopt the reconstruction loss $\mathcal{L}_{flow}$ to optimize the two embeddings (see Appendix A.3 for details).
3.3.4. Self-supervised region representation learning.
@@
Here, $\vec{h}^{k}_{r}$ ($k=1,2,3,4$) represents one of the four embeddings, and $\textsf{Decoder}^{k}$ is the corresponding decoder (implemented by a linear layer). This fusion module is easy to be extended to more views. The final loss function can be formulated as follows:
@@
4. Prompt with Task-specific Knowledge
We propose two graph-based prompt tuning methods, a manually-designed prompt method and a task-learnable prompt learning method, to extract explicit and implicit task-specific knowledge.
@@ 4.1. Manually-designed Prompt
In the manually-designed prompt, we adjust the subgraph structure, i.e., defining the graph pattern (i.e., prompt) $\mathcal{P}(P_{V},P_{E},\theta)$ in Section 3.2, to emphasize specific relationships and entities relevant to the target task, as shown in Figure 4.
Figure 4. Manually-designed prompt.
We enhance this process by manually adjusting the proportion of different types of edges and entities in the subgraph, thereby influencing the resultant embeddings. To achieve this, we first define a set of task-specific entity weights $\mathbf{W}_{task}$ that reflect the importance of various types of entities for the downstream task. These weights are designed based on the domain knowledge and specific requirements of task. Formally, we define $\mathbf{W}_{task}=\{w_{\tau_{i}}\mid\tau_{i}\in P_{V}\},$ where $P_{V}$ is the set of all possible node types in region-induced subgraph, and $w_{\tau_{i}}\in[0,1]$ represents the importance of entity type $\tau_{i}$. The weights $w_{\tau_{i}}$ are manually defined and used as the deletion proportion for each entity type. We define a function $\textsf{Adjust}(\cdot)$ to delete nodes based on $\mathbf{W}_{task}$$\textsf{Adjust}(g_{r},\mathbf{W}_{task})\rightarrow g^{\prime}_{r}$, where $g_{r}=(V_{r},E_{r})$ is the original subgraph. The adjusted subgraph $g^{\prime}_{r}=(V^{\prime}_{r},E^{\prime}_{r})$ is input to the pre-trained GURP to generate the task-specific embeddings. The function Adjust obtains the adjusted subgraph $g^{\prime}_{r}=(V^{\prime}_{r},E^{\prime}_{r})$ as follows:
@@
The manually-designed prompt is the graph pattern $\mathcal{P}$ with the set of task-specific entity weights $\mathbf{W}_{task}$, which guide the $\textsf{Adjust}(\cdot)$ function in modifying the subgraph $g_{r}$ to $g^{\prime}_{r}$. Incorporating such prompts, we can tailor the urban region representations to better capture task-relevant entities and relationships.
@@ 4.2. Task-learnable Prompt
Apart from leveraging explicitly designed task-specific knowledge prompts to enrich region representations, we can also learn prompts based on the task data and pre-trained representations to induce implicit knowledge within the task data, thereby enhancing region representations.
Considering that we describe the attributes of an urban region through subgraph induced with region nodes from $G$, it is intuitive to describe the task adaptation to different region attributes with graphs as well. Therefore, we propose a prompting module to use prompt graphs to portray the adaptation of downstream task relative to the original region embedding, as shown in Figure 5.
@@
We introduce the graph kernel method (Vishwanathan et al., 2010; Feng et al., 2022) to measure the similarity between the prompt graphs and the region subgraphs due to its performance and computational efficiency. Let $\mathcal{P}$ denote the set (i.e., prompt) consisting of $m$ trainable (hidden) prompt graphs, where each graph $p\in\mathcal{P}$ is defined as $p=(V^{p},E^{p},\mathbf{X}^{p})$. Here, $\mathbf{X}^{p}\in\mathbb{R}^{|V^{p}|\times d}$ denotes the node attribute matrix, where $d$ is the attribute dimension. For each region subgraph $g_{r}$, we utilize an attribute graph kernel function $K_{p}(\cdot,\cdot)$ to measure the similarity between $g_{r}$ and each prompt graph $p_{1},p_{2},\ldots,p_{m}$. Specifically, we employ the $P$-step random walk kernel (more details in Appendix A.4) to compute the kernel value between region subgraph $g_{r}$ and each prompt graph $p_{i}$. These values are concatenated to form an $m$-dimensional prompt vector $\vec{h}_{\mathcal{P}}\in\mathbb{R}^{m}$ for each region.
The prompt vector is used as a data operator to modify the pre-trained graph model. Conceptually, we can concatenate $\vec{h}_{\mathcal{P}}$ with each initial node feature vector $\vec{v}$ and feed the result into a linear layer $\textsf{Linear}(\cdot)$, such that each entity is modified by $\vec{v}_{*}=\textsf{Linear}(\vec{h}_{\mathcal{P}}||\vec{v})$. To reduce the complexity, in practice, we concatenate $\vec{h}_{\mathcal{P}}$ with the pre-trained region embedding $\vec{h}_{r}$ instead of each entity embedding. To learn the prompting model, we update the prompt graphs and linear layer based on specific downstream task with a simple MSE loss function.
Table 1. Main performance comparison of different models on downstream tasks. The best results for each metric are highlighted in bold, the second-best are underlined, and the third-best are marked with an asterisk.
Model New York City (NYC) Chicago (CHI)
Crime Check-in Carbon Income Crime Crash Carbon Education
MAE RMSE MAE RMSE MAE RMSE MAE RMSE MAE RMSE MAE RMSE MAE RMSE MAE RMSE
TransR-N 111.32 148.86 534.70 805.28 115.65 175.73 36799 45098 209.07 307.45 106.91 176.86 97.15 126.59 653.68 932.56
TransR-G 109.22 146.91 530.43 802.43 116.80 179.25 36363 44571 208.40 306.16 104.68 174.42 95.76* 124.37* 601.97 875.24
node2vec 138.46 175.80 573.12 837.44 174.45 225.42 45330 56182 246.07 349.97 121.90 174.91 115.62 148.95 832.89 1139.33
GAE 96.85 126.36 349.94 542.92 115.99 174.35 35338 43200 200.61* 300.24 100.54 173.32 95.29 123.66 702.82 986.70
MVURE 72.47* 99.84 310.97 498.08 99.30 151.01 24031 35351 206.08 289.86* 96.95 148.87* 98.19 130.31 593.80 843.10
MGFN 73.50 97.44* 300.15 456.63* 136.49 185.82 30266 40717 214.02 310.83 108.37 179.85 96.56 125.50 709.91 998.03
HRE 76.93 102.07 329.55 513.80 102.95 160.58 25903 35418 206.61 298.39 99.95 158.63 97.11 133.00 555.56 806.60
HREP 74.85 100.04 320.34 505.52 102.49 159.68 24643 34148 206.59 298.24 99.77 158.66 97.19 132.91 555.01 806.36
ReCP 73.85 100.56 292.87* 524.89 95.55 150.84 24440* 35044 207.49 298.97 93.30* 157.52 95.84 125.73 455.98* 708.37*
GURP 72.09 93.22 271.69 412.92 108.17 158.03 25268 34943* 161.27 225.24 80.36 108.78 97.83 133.03 455.43 677.36
GURPP 71.48 91.61 261.45 388.23 100.53* 152.70* 22902 31945 146.39 203.05 65.30 92.11 64.844 88.56 420.62 656.98
Enhance 1.36% 5.98% 10.73% 14.98% -4.95% -1.21% 4.70% 6.45% 27.03% 29.95% 30.01% 38.12% 31.95% 28.39% 7.76% 7.26%
@@ 5. Experiments
We conduct extensive experiments on six downstream tasks of two cities from the following perspectives: (1) overall performance comparison with SOTA baselines, (2) effectiveness analysis of two prompting methods, (3) few-shot and zero-shot generalization capabilities, (4) ablation studies, (5) scalability evaluation, and (6) comparison with language-based prompting methods.
@@ 5.1. Experimental Setup
We collect the datasets of region division, POI, urban imagery, taxi trips, road network and six downstream tasks for two representative cities in the United States: New York City (NYC) and Chicago (CHI).
@@
We compare our model with the following representative baselines: knowledge graph embedding methods and its variants (TransR (Lin et al., 2015), TransR-N, TransR-G), graph embedding methods (node2vec (Grover and Leskovec, 2016), GAE (Kipf and Welling, 2016)), and urban region embedding methods (MVURE (Zhang et al., 2021), MGFN (Wu et al., 2022), HREP (Zhou et al., 2023a),ReCP (Li et al., 2024a)). We follow the descriptions in these works to preprocess the data into the corresponding format and run the baselines with the parameter settings specified in the papers to ensure fairness.
@@
We use three widely recognized evaluation metrics to evaluate the prediction performance: the coefficient of determination ($R^{2}$), root mean squared error (RMSE), and mean absolute error (MAE). Better performance is indicated by a higher $R^{2}$ and lower RMSE and MAE values. Our model includes three version: 1) GURP represents the pre-train module, 2) $\textbf{GURPP}_{M}$ is GURP with manually-designed prompt, and 3) $\textbf{GURPP}_{T}$ is GURP with task-learnable prompt. Details on datasets and implementation are in Appendix A.5.
5.2. Overall Performance
@@
We evaluate the overall performance of the baselines and our model on datasets collected from two cities, covering six different downstream tasks. The results of the experiments are shown in Table 1, from which we find the following.
(1) GURPP demonstrates superior performance across both cities and tasks. In the pre-training phase, GURP achieves state-of-the-art performance on 6 out of 8 tasks, with particularly significant improvements on CHI datasets (27.03%-31.95% MAE reduction on Crime/Crash prediction). When enhanced with prompting methods, GURPP further improves performance by an average of 11.30% (MAE) and 10.16% (RMSE) across all tasks compared to GURP. The consistent improvements across heterogeneous tasks (12.46% average MAE gain and 14.62% RMSE gain over baselines) validate the effectiveness of our pre-training and prompting framework.
@@
(2) The results highlight our model’s exceptional adaptability to different urban contexts. Specifically, GURPP achieves average improvements of 24.2% (MAE) and 25.9% (RMSE) in CHI versus 3.0% (MAE) and 6.6% (RMSE) in NYC. This may be due to the fact that the baselines tend to be effective only at extracting information specific to certain data types, but they do not consider spatial structure and functional layouts among urban entities. For example, existing baselines rely on the flow view for information extraction, but the sparse content in the CHI dataset significantly degrades downstream task performance. In contrast, GURPP performs excellently on both urban datasets, demonstrating its robustness and universality across diverse urban environments.
@@
(3) To further validate the effectiveness of our subgraph-centric and multi-view fusion module, we conduct experiments comparing the performance of models with and without these modules. Specifically, the result that TransR-G outperforms TransR-N by an average of 1.27% (MAE) indicates that the subgraph-centric method can preserve the structural and contextual information of regions, providing more effective urban region embeddings. The average 32.55% (MAE) improvement for GURP over TransR-G is mainly due to the heterogeneous feature extraction, which improves the model’s comprehensive understanding of urban features.
5.3. Effectivness of Prompting Methods
5.3.1. Manually-designed Prompt Study.
In the real world, crime rates are closely linked to specific POI categories, roads, and junctions, with locations like ATMs, bars, and parks being more crime-prone. Road density and junction complexity also impact crime. Regarding user check-ins, brand popularity plays a key role, with users often preferring well-known brands like Starbucks or McDonald’s. Based on these observations, we design two patterns: P1 (ignoring POI brand information) and P2 (randomly sampling POI, road, and junction nodes with a 0.9 ratio) to evaluate the performance of manually-designed prompts across different tasks.
Table 2. Performance of manually-designed prompt $\text{GURPP}_{M}$.